OpenStudy (anonymous):

You can use the Law of Sines if you are given a = 31, b = 52 and m

OpenStudy (aravindg):

\[\large \dfrac{a}{\sin A}=\dfrac{b}{\sin B}=\dfrac{c}{\sin C}\]

OpenStudy (aravindg):

^^use that

OpenStudy (anonymous):